OpenStudy (anonymous):

Algebra 2 please help!!!!

OpenStudy (anonymous):

n^4-11n^2+30/n^4-7n^2+10 simplify

OpenStudy (amistre64):

you have to factor the top and bottom to see what like factors can cancel out

OpenStudy (anonymous):

i think i got the answer hold on just a minute if u dont ,mind

OpenStudy (anonymous):

i don't know how to factor it.... and ok :)

OpenStudy (amistre64):

is it becasue of the ^4 and ^2? or just becasue you dont know how to factor?

OpenStudy (anonymous):

yeah cuz of the ^4 and ^2

OpenStudy (amistre64):

its just a disquised quadratic: n^4 = (n^2)^2 so we can rewrite this cleaner by letting x = n^2

OpenStudy (amistre64):

n^4-11n^2+30 -------------- n^4-7n^2+10 (n^2)^2-11n^2+30 ----------------- (n^2)^2-7n^2+10 x^2-11x+30 ------------ x^2-7x+10

OpenStudy (amistre64):

now you should be able to recognize it as a more normal quadratic looking setup

OpenStudy (anonymous):

yeah so would it be.... (x -5)(x-6)/(x-2)(x-5)

OpenStudy (amistre64):

correct, and the x-5 parts can cancel

OpenStudy (anonymous):

so x-6/x-2...? are there any restirctions?

OpenStudy (anonymous):

oh wait those should be n^2 iinstead of x right?

OpenStudy (amistre64):

x not equal 2 fer sure; but you should prolly want to include the original restriction with x=5 as out. recall that we said let n^2 = x therefore; x = sqrt(n) and we can continue the simplifications with that\[\frac{\sqrt n-6}{\sqrt n-2}\]

OpenStudy (amistre64):

depending on the textbook you have to work from, i believe that most simplifications in general hate sqrts in the bottom, so we would have to run a conjugate on it

OpenStudy (anonymous):

they just want it like n^2-6/n^2-2 where n does not equal 5 or 2... :) thanks for all your help!!

OpenStudy (amistre64):

if there are any options to guide us towards, that would be nice :)

OpenStudy (amistre64):

cool

OpenStudy (amistre64):

n not equal to +-sqrt(2) or +-sqrt(5) would be better restrictions :)

OpenStudy (amistre64):

lol, i see my mismath in that n=x^2; x = sqrt n post ... good eye